OpenStudy (anonymous):

Find the average rate of change for the following functions on the indicated intervals. f(x)= xcubed - 2x; (0,4) just don't understand how to work the problem.Help?

OpenStudy (saifoo.khan):

HINT: rate of change is it's derivative.

jimthompson5910 (jim_thompson5910):

The average rate of change from 0 to 4 is given by ( f(4) - f(0) )/( 4 - 0 )

OpenStudy (anonymous):

still dont understand

jimthompson5910 (jim_thompson5910):

saifoo.khan, that's the instantaneous rate of change

jimthompson5910 (jim_thompson5910):

First find f(4), do you know how to do that?

OpenStudy (anonymous):

don't remember

OpenStudy (saifoo.khan):

Ah, then what's this called? @jim_thompson5910 What's the main difference? Thanks for correcting.

jimthompson5910 (jim_thompson5910):

f(x)= x^3 - 2x f(4)= (4)^3 - 2*(4) <<--- Evaluate this to find f(4)

OpenStudy (anonymous):

but 4 is the y

jimthompson5910 (jim_thompson5910):

No, 4 is the x

jimthompson5910 (jim_thompson5910):

f(4) means find f(x) when x = 4

OpenStudy (campbell_st):

For average rate of change... its the gradient of the secant across the interval x = 0 to x = 4 start by finding f(0) and f(4) then the average rate of change is \[\frac{f(4) -f(0)}{4 - 0}\]

OpenStudy (anonymous):

okay so find the function of 4 and 0?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (saifoo.khan):

So this is basically the average? @jim_thompson5910

OpenStudy (anonymous):

what now?

jimthompson5910 (jim_thompson5910):

So what did you get for f(4)?

OpenStudy (anonymous):

40

jimthompson5910 (jim_thompson5910):

Try it again, f(4) is NOT 40

OpenStudy (anonymous):

56

jimthompson5910 (jim_thompson5910):

it's the the average rate of change as stated at the top of the problem saifoo.khan

jimthompson5910 (jim_thompson5910):

good, you got it

jimthompson5910 (jim_thompson5910):

so f(4) = 56 what about f(0)?

OpenStudy (anonymous):

thats 0

jimthompson5910 (jim_thompson5910):

right again

jimthompson5910 (jim_thompson5910):

so ( f(4) - f(0) )/( 4 - 0 ) then becomes ( 56 - 0 )/( 4 - 0 )

jimthompson5910 (jim_thompson5910):

evaluate/simplify that to get your final answer

OpenStudy (anonymous):

so 14?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thats the average rate of change?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

think of it as the slope of the line that connects the two points (0,0) and (4,56)

OpenStudy (anonymous):

you good at word problems?

jimthompson5910 (jim_thompson5910):

sure, what's the question?

OpenStudy (anonymous):

same directions. a car travels 360 miles in a period of 180 minutes. find the average velocity of the car in miles per hour over this time period. isn't there a formula for this?

jimthompson5910 (jim_thompson5910):

first convert 180 minutes to hours

jimthompson5910 (jim_thompson5910):

Use the idea that 1 hour = 60 minutes

OpenStudy (anonymous):

3 hours

jimthompson5910 (jim_thompson5910):

So the car travels 360 miles in 3 hours The average velocity is then (Distance)/(Time)

OpenStudy (anonymous):

so 120?

jimthompson5910 (jim_thompson5910):

yes, the car's average velocity is then 120 mi/hr

OpenStudy (anonymous):

thanks so much! i have lots more to do..are you busy?

jimthompson5910 (jim_thompson5910):

no not really, which ones are bugging you the most?

OpenStudy (anonymous):

natural logarithms..just don't remember them

jimthompson5910 (jim_thompson5910):

A natural log is simply a logarithm, but it has a special base known as 'e'

jimthompson5910 (jim_thompson5910):

The value of 'e' is roughly e = 2.718...

OpenStudy (anonymous):

5/e^x +1 =1 solve for x

OpenStudy (anonymous):

thats whats bugging me..i can't simplify it far enough

jimthompson5910 (jim_thompson5910):

\[\Large \frac{5}{e^x}+1=1\] or \[\Large \frac{5}{e^x+1}=1\] ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

which one?

OpenStudy (anonymous):

the second

jimthompson5910 (jim_thompson5910):

Start by multiplying both sides by e^x+1 to get \[\Large 5=1(e^x+1)\] \[\Large 5=e^x+1\] Is this one of your steps?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

subtract 1

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

to get what?

OpenStudy (anonymous):

but then im stuck

OpenStudy (anonymous):

i have 4= e^x

jimthompson5910 (jim_thompson5910):

ok great so far

OpenStudy (anonymous):

but now what?

jimthompson5910 (jim_thompson5910):

now take the natural log of both sides

OpenStudy (anonymous):

1.39

OpenStudy (anonymous):

e cancels right?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

if the book want's an exact answer, then simply leave it as x = ln(4)

jimthompson5910 (jim_thompson5910):

but if it wants an approximate answer, then you would write x = 1.39

OpenStudy (anonymous):

okay..well..it's a summer assignment..i should be okay just as long as i show work

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

\[e ^{2x}+2+e ^{-2x}\]

OpenStudy (anonymous):

help?

jimthompson5910 (jim_thompson5910):

simplifying this? or solving?

OpenStudy (anonymous):

factoring

OpenStudy (anonymous):

into binomials

jimthompson5910 (jim_thompson5910):

hmm one sec

jimthompson5910 (jim_thompson5910):

ah ok I think I see what to do

jimthompson5910 (jim_thompson5910):

\[e^{2x}+2+e^{-2x}\] is the same as \[e ^{2x}+2+\frac{1}{e^{2x}}\]

OpenStudy (anonymous):

yeah..just an inverse

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

so they cancel?

OpenStudy (anonymous):

or make 1

jimthompson5910 (jim_thompson5910):

Not quite

OpenStudy (anonymous):

common denomenator?

jimthompson5910 (jim_thompson5910):

Yes, you need to find the LCD

jimthompson5910 (jim_thompson5910):

then get each term to have that LCD, then combine the fractions

OpenStudy (anonymous):

so it would be e to the 2x underneath and one fraction would have e^2x squared on top

jimthompson5910 (jim_thompson5910):

yes, that's the first fraction

OpenStudy (anonymous):

okay..now what

jimthompson5910 (jim_thompson5910):

so \[e ^{2x}+2+\frac{1}{e^{2x}}\] becomes \[\frac{\left(e^{2x}\right)^2}{e^{2x}}+2+\frac{1}{e^{2x}}\]

jimthompson5910 (jim_thompson5910):

Do the same with the middle term 2

OpenStudy (anonymous):

put it over e?

jimthompson5910 (jim_thompson5910):

well multiply it by \[\Large \frac{e^{2x}}{e^{2x}}\]

OpenStudy (anonymous):

okay.. now what

jimthompson5910 (jim_thompson5910):

\[\frac{\left(e^{2x}\right)^2}{e^{2x}}+2+\frac{1}{e^{2x}}\] then becomes \[\frac{\left(e^{2x}\right)^2}{e^{2x}}+\frac{2e^{2x}}{e^{2x}}+\frac{1}{e^{2x}}\]

jimthompson5910 (jim_thompson5910):

Now combine the fractions

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

combined